OpenStudy (anonymous):

The coefficient of linear expansion of iron is 1.0 x 10-5 per oC. The surface area of an iron cube, with an edge length of 5.0 cm, will increase by what amount if it is heated from 10 oC to 60 oC?

OpenStudy (anonymous):

i get 2.5*10^-3 and that is wrong :(

OpenStudy (anonymous):

the formula is ΔL=LαΔT

OpenStudy (ujjwal):

coefficient of superficial expansion, \(\beta\)=2\(\alpha\) Now use relation \(\Delta A=A\beta\Delta T\) to find increase in area.

OpenStudy (anonymous):

ohhh crap

OpenStudy (anonymous):

totally forgot about volume expansion derrr thanks

OpenStudy (anonymous):

β=3α by the way

OpenStudy (ujjwal):

that's not volume expansion. that's area expansion. \(\beta=2\alpha\) and \(\gamma=3\alpha\) we use \(\gamma\) for volumetric expansion.

OpenStudy (ujjwal):

use \(\beta\) if its expansion for area.

OpenStudy (anonymous):

ah k i will just try it..gimme a sec

OpenStudy (ujjwal):

you asked for a sec, I will give you ten.. :D

OpenStudy (anonymous):

ok so i still get the wrong answer...

OpenStudy (anonymous):

i did 2*1*10-5*50*5^2=5*10^-3

OpenStudy (anonymous):

2*1*10^-5*50*5^2=5*10^-3

OpenStudy (anonymous):

(2)*(1*10-5)*(50)*(5^2)=5*10^-3 i think the 5^2 is wrong?

OpenStudy (ujjwal):

use meter not centimeter for length, and then calculate area.

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

what do you get? i know the answer but yeh

OpenStudy (ujjwal):

calculate properly, you will get 2.5*10^-6 m^2 . If the answer is in cm^2, just convert it to cm^2

OpenStudy (ujjwal):

multiply it by 1/2??

OpenStudy (anonymous):

lol but thats the wrong answer. it is 0.15cm^2

OpenStudy (ujjwal):

The question is asking for total area at last and i calculated the difference. so you need to add it to original area. As simple as that.

OpenStudy (anonymous):

so total change is 0.00025cm^2 and originally it is 25cm^2 but that doesn't equal 0.15cm^2 if you find the difference

sam (.sam.):

you two asking the same question? http://openstudy.com/study?login#/updates/4fd21a4ee4b057e7d2211b72

OpenStudy (anonymous):

lol yeh looks like it

OpenStudy (anonymous):

no i don't have two accounts

OpenStudy (anonymous):

and no i don't know that person